# ZIP-RC: Optimizing Test-Time Compute via Zero-Overhead Joint Reward-Cost Prediction

### v2

@@ 1 Introduction
A promising approach is the canonical test-time scaling method Best-of-N (BoN) sampling, which generates NN candidates and selects the best using a learned verifier, reward model, or majority vote (cobbe2021training; zheng2023judging; kwon2023reward; lightman2023let; wang2022self). While appealing in theory due to its parallelism, BoN is not adaptive: every trajectory is carried to completion regardless of promise. On easy tasks this wastes computation, and on hard tasks it inflates latency, since wall-clock time is governed by the longest generation and both length and total compute grow with NN (leviathan2023fast). What is missing is a way for models to anticipate which samples are worth continuing and which should be paused or abandoned, so that parallel effort is concentrated on trajectories most likely to succeed and fastest to complete.
@@ 3 Preliminaries
Best-of-NN (BoN) is an inference-time selection mechanism that decouples generation from evaluation to improve output quality. Given a prompt $\mathbf{x}$ and a generator policy $\pi$, the method draws NN independent and identically distributed (i.i.d.) terminated states $s_{T}^{(1)},\dots,s_{T}^{(N)}$ from the policy. A learned verifier $\hat{V}:\mathcal{V}^{*}\to\mathbb{R}$, typically a reward model, then assigns a scalar score to each terminated state. The final output is the state with the highest score, selected as
@@ 4 Zero-Overhead Inference-time Prediction of Reward & Cost
$\mathcal{L}(s_{t})=\mathcal{L}_{\text{aux}}(s_{t})+\alpha_{\mathrm{KL}}\;\mathrm{KL}\!\left(\pi_{\theta}(\cdot\mid s_{t})\;\middle\|\;\pi(\cdot\mid s_{t})\right).$ (3)
@@ 5 Test-time Compute using ZIP-RC (ZIP-RC sampling)
Existing test-time methods, however, are heuristic and often inefficient. BoN can, in principle, explore as much as all other approaches with a large enough NN, but this is impractical given compute and latency constraints. While more sophisticated search strategies like beam search try to explore more efficiently by allowing for intermediate branching and pruning at intervals, one could imagine removing constraints on search further. Though the goal of test-time search is clear—maximize task success while minimizing generation cost—prior methods do not achieve this in a principled way. Our goal is to propose a method that does. In this section, we introduce ZIP-RC sampling, which leverages predictions from ZIP-RC to explicitly optimize generations for both success and cost.
@@
A strategy $\mu$ is a policy in this meta-MDP: at each timestep it maps the current prefix tree $S_{t}$ to a multiset $A_{t}=\mu(S_{t})$ of nodes to expand, and at horizon $T$ selects a finished $f^{*}$. BoN corresponds to placing NN copies of the root in $A_{0}$ and thereafter always expanding every unfinished leaf until completion, finally selecting the highest-scoring candidate. Beam search with width $B$ instead enforces $|A_{t}|=B$ at all times: at most steps $A_{t}$ is just the $B$ current leaves, but at pruning intervals of length $k$ it ranks leaves by a score, discards the weakest $p$, and duplicates stronger ones so that the frontier is refilled back to $B$, thereby pruning and branching in a controlled manner before ultimately selecting the best finished prefix at $T$.
@@ 5.2 Computing an Optimal Zero-overhead Search Strategy
We can observe that the expression $Q^{\mathrm{Rollouts}}$ contains several interpretable quantities. Namely, the expected maximum value quantifies the marginal benefit of branching or pruning; the incremental gain from increasing NN is large when the value distribution has high variance, and conversely, the marginal loss of pruning is small under low variance. Furthermore, the expected maximum remaining tokens and the expected total tokens capture the marginal cost of branching; increasing NN always increases the expected total remaining tokens and the maximum remaining length, which drives up latency, and pruning will always reduce the cost.
@@ 6.1 Experimental Setup
Baselines. We evaluate against the following baselines that consist of popular sampling strategies that fall under the parallel sampling paradigm where multiple candidate samples are generated in parallel and there is some selection method. Other notable paradigms include beam search or self-refinement. However, we use parallel sampling methods, which are the most commonly used and reported as they do not suffer from collapsing diversity issues that arise from branching and generating with similar prefixes or the ballooning latency issues from methods that generate samples sequentially. We use stronger adaptations of Best-of-NN (BoNN), and an ablation of ZIP-RC that performs pruning without the sampling utility optimization and instead uses the expected reward directly.:
@@
Metrics. First and foremost we measure accuracy on each benchmark as it is an obvious and good measure for performance and high-quality responses. Beyond performance, we measure efficiency and latency. Normalized compute reports total FLOPs per prompt normalized by the FLOPs of a single-sample generation for that prompt. We compute FLOPs with the standard $2N$ rule (proportional to the sum of input and generated tokens) and account for KV caching where applicable. Normalized best-case latency measures the lower bound on wall-clock time as the maximum number of sequential forward passes across the candidate set; with unconstrained data-parallel sampling, latency is governed by the longest trajectory. Generation cost aggregates these via a linear combination, $\text{GenCost}=\alpha\cdot\text{NormCompute}+(1-\alpha)\cdot\text{NormLatency}$. Unless otherwise specified, we use $\alpha=0.1$, which roughly balances compute and latency in typical parallel regimes (e.g., eight parallel samples often behave like two to three serial generations in practice). For ZIP-RC sampling we sweep $\beta$, which trades off expected quality against cost in the utility; when reporting matched-cost comparisons we set $\beta=0.005$ and cap the pool at 88 samples for fair comparison to other baselines.
@@ 6.3 Tracing the quality–compute–latency frontier
We next test whether maximizing the sampling utility with specific cost coefficients achieves controllable tradeoffs. At each decision point, ZIP-RC evaluates meta-actions that serve three complementary purposes. First, initiating new samples only when necessary and avoiding continuing to generate low-value trajectories saves compute, which is reflected in the compute bound setting in the bottom half of  fig. 3 ($\alpha=1.0$), where ZIP-RC achieves compute savings. Second, penalizing the continued sampling of long outliers avoids samples that would dominate latency. Third, expanding the initial pool of samples while planning to use a near-term maximum horizon enables the search to pursue early finishers without paying the full wall-clock cost of long runs. These mechanisms together drive the latency savings observed in the top half of  fig. 3 ($\alpha=0.0$). In both settings $\beta$ is successfully used similar to NN in BoN in order to increase performance for more generation cost. Parameters $\alpha$ and $\beta$ together thus provide simple control knobs over compute–latency emphasis and quality–cost trade-off.
@@ 6.4 Adaptive inference with ZIP-RC sampling
At matched cost, ZIP-RC sampling improves accuracy over MV and weighted BoNN on all models and benchmarks. On harder subsets such as AIME 2024, gains reach up to 12% absolute while using less average cost. The adaptive policy naturally uses more samples when the predicted reward distribution is high-variance—where the expected benefit of best-of-NN is greatest—and conserves compute when one trajectory is expected to be dominant. This pattern is evident on the mixed-difficulty benchmark (left-most column in  fig. 3) and across model scales: weaker models and harder tasks receive more samples, leading to higher overall accuracy.
Takeaways. ZIP-RC’s real-time predictions are accurate and reliable enough to enable principled search during decoding. This yields (i) reliable mid-generation detection of weak or overlong trajectories, (ii) smooth and tunable Pareto frontiers between quality, compute, and latency, and (iii) adaptive allocation that consistently outperforms fixed-budget Best-of-NN at the same or lower cost.

### v3

# Zero-Overhead Introspection for Adaptive Test-Time Compute

@@ 1 Introduction
A promising approach is the canonical test-time scaling method Best-of-N (BoN) sampling, which generates $N$ candidates and selects the best using a learned verifier, reward model, or majority vote (cobbe2021training; zheng2023judging; kwon2023reward; lightman2023let; wang2022self). While appealing in theory due to its parallelism, BoN is not adaptive: every trajectory is carried to completion regardless of promise. On easy tasks this wastes computation, and on hard tasks it inflates latency, since wall-clock time is governed by the longest generation and both length and total compute grow with $N$ (leviathan2023fast). What is missing is a way for models to anticipate which samples are worth continuing and which should be paused or abandoned, so that parallel effort is concentrated on trajectories most likely to succeed and fastest to complete.
@@ 3 Preliminaries
Best-of-NN (BoN) is an inference-time selection mechanism that decouples generation from evaluation to improve output quality. Given a prompt $\mathbf{x}$ and a generator policy $\pi$, the method draws $N$ independent and identically distributed (i.i.d.) terminated states $s_{T}^{(1)},\dots,s_{T}^{(N)}$ from the policy. A learned verifier $\hat{V}:\mathcal{V}^{*}\to\mathbb{R}$, typically a reward model, then assigns a scalar score to each terminated state. The final output is the state with the highest score, selected as
@@ 4 Zero-Overhead Inference-time Prediction of Reward & Cost
$\mathcal{L}(s_{t})=\mathcal{L}_{\text{aux}}(s_{t})+\alpha_{\mathrm{KL}}\;\mathrm{KL}\!\left(\pi(\cdot\mid s_{t})\;\middle\|\;\pi_{\theta}(\cdot\mid s_{t})\right).$ (3)
@@ 5 Test-time Compute using ZIP-RC (ZIP-RC sampling)
Existing test-time methods, however, are heuristic and often inefficient. BoN can, in principle, explore as much as all other approaches with a large enough $N$, but this is impractical given compute and latency constraints. While more sophisticated search strategies like beam search try to explore more efficiently by allowing for intermediate branching and pruning at intervals, one could imagine removing constraints on search further. Though the goal of test-time search is clear—maximize task success while minimizing generation cost—prior methods do not achieve this in a principled way. Our goal is to propose a method that does. In this section, we introduce ZIP-RC sampling, which leverages predictions from ZIP-RC to explicitly optimize generations for both success and cost.
@@
A strategy $\mu$ is a policy in this meta-MDP: at each timestep it maps the current prefix tree $S_{t}$ to a multiset $A_{t}=\mu(S_{t})$ of nodes to expand, and at horizon $T$ selects a finished $f^{*}$. BoN corresponds to placing $N$ copies of the root in $A_{0}$ and thereafter always expanding every unfinished leaf until completion, finally selecting the highest-scoring candidate. Beam search with width $B$ instead enforces $|A_{t}|=B$ at all times: at most steps $A_{t}$ is just the $B$ current leaves, but at pruning intervals of length $k$ it ranks leaves by a score, discards the weakest $p$, and duplicates stronger ones so that the frontier is refilled back to $B$, thereby pruning and branching in a controlled manner before ultimately selecting the best finished prefix at $T$.
@@ 5.2 Computing an Optimal Zero-overhead Search Strategy
We can observe that the expression $Q^{\mathrm{Rollouts}}$ contains several interpretable quantities. Namely, the expected maximum value quantifies the marginal benefit of branching or pruning; the incremental gain from increasing $N$ is large when the value distribution has high variance, and conversely, the marginal loss of pruning is small under low variance. Furthermore, the expected maximum remaining tokens and the expected total tokens capture the marginal cost of branching; increasing $N$ always increases the expected total remaining tokens and the maximum remaining length, which drives up latency, and pruning will always reduce the cost.
@@ 6.1 Experimental Setup
Baselines. We evaluate against the following baselines that consist of popular sampling strategies that fall under the parallel sampling paradigm where multiple candidate samples are generated in parallel and there is some selection method. Other notable paradigms include beam search or self-refinement. However, we use parallel sampling methods, which are the most commonly used and reported as they do not suffer from collapsing diversity issues that arise from branching and generating with similar prefixes or the ballooning latency issues from methods that generate samples sequentially. We use stronger adaptations of Best-of-$N$ (BoNN), and an ablation of ZIP-RC that performs pruning without the sampling utility optimization and instead uses the expected reward directly.:
@@
Metrics. First and foremost we measure accuracy on each benchmark as it is an obvious and good measure for performance and high-quality responses. Beyond performance, we measure efficiency and latency. Normalized compute reports total FLOPs per prompt normalized by the FLOPs of a single-sample generation for that prompt. We compute FLOPs with the standard $2N$ rule (proportional to the sum of input and generated tokens) and account for KV caching where applicable. Normalized best-case latency measures the lower bound on wall-clock time as the maximum number of sequential forward passes across the candidate set; with unconstrained data-parallel sampling, latency is governed by the longest trajectory. Generation cost aggregates these via a linear combination, $\text{GenCost}=\alpha\cdot\text{NormCompute}+(1-\alpha)\cdot\text{NormLatency}$. Unless otherwise specified, we use $\alpha=0.1$, which roughly balances compute and latency in typical parallel regimes (e.g., eight parallel samples often behave like two to three serial generations in practice). For ZIP-RC sampling we sweep $\beta$, which trades off expected quality against cost in the utility; when reporting matched-cost comparisons we set $\beta=0.005$ and cap the pool at $8$ samples for fair comparison to other baselines.
@@ 6.3 Tracing the quality–compute–latency frontier
We next test whether maximizing the sampling utility with specific cost coefficients achieves controllable tradeoffs. At each decision point, ZIP-RC evaluates meta-actions that serve three complementary purposes. First, initiating new samples only when necessary and avoiding continuing to generate low-value trajectories saves compute, which is reflected in the compute bound setting in the bottom half of  fig. 3 ($\alpha=1.0$), where ZIP-RC achieves compute savings. Second, penalizing the continued sampling of long outliers avoids samples that would dominate latency. Third, expanding the initial pool of samples while planning to use a near-term maximum horizon enables the search to pursue early finishers without paying the full wall-clock cost of long runs. These mechanisms together drive the latency savings observed in the top half of  fig. 3 ($\alpha=0.0$). In both settings $\beta$ is successfully used similar to $N$ in BoN in order to increase performance for more generation cost. Parameters $\alpha$ and $\beta$ together thus provide simple control knobs over compute–latency emphasis and quality–cost trade-off.
@@ 6.4 Adaptive inference with ZIP-RC sampling
At matched cost, ZIP-RC sampling improves accuracy over MV and weighted BoNN on all models and benchmarks. On harder subsets such as AIME 2024, gains reach up to 12% absolute while using less average cost. The adaptive policy naturally uses more samples when the predicted reward distribution is high-variance—where the expected benefit of best-of-$N$ is greatest—and conserves compute when one trajectory is expected to be dominant. This pattern is evident on the mixed-difficulty benchmark (left-most column in  fig. 3) and across model scales: weaker models and harder tasks receive more samples, leading to higher overall accuracy.
Takeaways. ZIP-RC’s real-time predictions are accurate and reliable enough to enable principled search during decoding. This yields (i) reliable mid-generation detection of weak or overlong trajectories, (ii) smooth and tunable Pareto frontiers between quality, compute, and latency, and (iii) adaptive allocation that consistently outperforms fixed-budget Best-of-$N$ at the same or lower cost.